please help me solve this
z+7/9+1 = 2z+1/6

Answer by checkley75(3666)

Z+7/9+1=2Z+1/6
Z-2Z=1/6-7/8-1
-3Z=(4-3*7-24)/24
-3Z=(4-21-24)/24
-3Z=-41/24
Z=-41/24*-1/3
Z=41/72
----------------------------------
(Z+7)/9+1=(2Z+1)/6 YOU NEED TO USE ( ) AROUND THE COMPLEX NUMERATORS.
(Z+7+9)/9=(2Z+1)/6
(Z+16)/9=(2Z+1)/6 NOW CROSS MULTIPLY
9(2Z+1)=6(Z+16)
18Z+9=6Z+96
18Z-6Z=96-9
12Z=87
Z=87/12
Z=7.25 ANSWER
HOPE THIS IS THE RIGHT ANSWER!!!!!!!
USE ( ) WHEN IN DOUBT.